OpenStudy (anonymous):

Which of the following is a pair of equivalent fractions? ten-twelfths and thirty forty-eighths fourteen twenty-eighths and twenty-one thirty-fifths six twentieths and four-fifteenths seven-eighths and twenty-eight thirty-seconds

OpenStudy (diyadiya):

What you have to do is to simply the two fractions Lets do the first one.. \[\frac{10}{12}= \frac{5}{6}\]\[\frac{30}{48}= \frac{5}{8}\] therefore \[\frac{10}{12}\neq \frac{30}{48}\]

OpenStudy (diyadiya):

Lets check the second one \[\frac{14}{28}= \frac{1}{2}\]\[\frac{21}{35}= \frac{3}{5}\] Therefore \[\frac{14}{28}\neq \frac{21}{35}\]

OpenStudy (diyadiya):

Can you check the third one ?@kiddpolo43

OpenStudy (diyadiya):

@kiddpolo43 all you have to do is to simplify the fraction and check if they are equal or not

OpenStudy (anonymous):

okay

OpenStudy (diyadiya):

\[\frac{6}{20}= -\]\[\frac{4}{15}=-\] What do you get and are they equal ?

OpenStudy (anonymous):

3/10 2/7

OpenStudy (diyadiya):

3/10 is right But can you simplify 4/15 ?

OpenStudy (anonymous):

no not really

OpenStudy (diyadiya):

So 4/15 remain 4/15 so are they equal?

OpenStudy (anonymous):

nope

OpenStudy (diyadiya):

Good :) So now we can move on to the next and the last one . and check if they are equal or not

OpenStudy (anonymous):

okay

OpenStudy (diyadiya):

\[\frac{7}{8}= -\]\[\frac{28}{32}=-\] can you do it ?

OpenStudy (diyadiya):

@kiddpolo43 what do you get when you simplify 28/32 ?

OpenStudy (anonymous):

both are equal 7/8

OpenStudy (diyadiya):

Yes so ,seven-eighths and twenty-eight thirty-seconds is the pair of equivalent fractions :)

OpenStudy (anonymous):

yes